# Mitigating Hallucinations in YOLO-based Object Detection Models: A Revisit to Out-of-Distribution Detection

### v2

@@ Abstract
Object detection systems must reliably perceive objects of interest without being overly confident to ensure safe decision-making in dynamic environments. Filtering techniques based on out-of-distribution (OoD) detection are commonly added as an extra safeguard to filter hallucinations caused by overconfidence in novel objects. Nevertheless, evaluating YOLO-family detectors and their filters under existing OoD benchmarks often leads to unsatisfactory performance. This paper studies the underlying reasons for performance bottlenecks and proposes a methodology to improve performance fundamentally. Our first contribution is a calibration of all existing evaluation results: Although images in existing OoD benchmark datasets are claimed not to have objects within in-distribution (ID) classes (i.e., categories defined in the training dataset), around $13\%$ of objects detected by the object detector are actually ID objects. Dually, the ID dataset containing OoD objects can also negatively impact the decision boundary of filters. These ultimately lead to a significantly imprecise performance estimation. Our second contribution is to consider the task of hallucination reduction as a joint pipeline of detectors and filters. By developing a methodology to carefully synthesize an OoD dataset that semantically resembles the objects to be detected, and using the crafted OoD dataset in the fine-tuning of YOLO detectors to suppress the objectness score, we achieve a $88\%$ reduction in overall hallucination error with a combined fine-tuned detection and filtering system on the self-driving benchmark BDD-100K. Our code and dataset are available at: https://gricad-gitlab.univ-grenoble-alpes.fr/dnn-safety/m-hood.
@@ I INTRODUCTION
In image classification, research has shown (e.g. [1]) that deep neural network-based classifiers can be overly confident in their predictions. Since object detection involves both localization and classification, overconfidence also manifests itself. For dynamic driving tasks, it is crucial to eliminate the impact of overconfidence. Failure to do so can lead to unsafe scenarios, such as the unintended activation of automatic emergency braking due to falsely detected pedestrians.
Out-of-distribution (OoD) detection is one promising approach to mitigating the impact of overconfidence in new or unfamiliar situations. Within image classification, OoD detection techniques can filter anomalies and correct the result from the predicted class to the artificially introduced “OoD” class. Transferring the techniques from classification into object detection, techniques such as VOS [2], SAFE [3] and BAM [4] are implemented on top of Faster-RCNN [5] as an extra component to subsequently filter the detected and classified objects by setting these objects to be of class “OoD”. A recent study [6] migrates state-of-the-art OoD detection methods in classification to YOLO series of object detectors. Despite the superior performance of OoD detectors operated on Faster-RCNN, the performance is still far from acceptable and reveals a significant improvement potential in YOLO.
@@
In this paper, we analyze the reasons behind the unsatisfactory OoD detection performance and design a methodology to improve the overall performance of mitigating hallucination errors (i.e., false detections) when encountering OoD samples. Our first result is identifying dataset quality issues and benchmark design in evaluating OoD used in VOS [2] and all other techniques, contributing to the high false positive rates observed in previous studies. Specifically, we uncover two critical problems and study their impact both in theory and concrete empirical evaluations:
@@
To further reduce the hallucination (overconfident predictions) caused by OoD samples, we introduce a novel approach to forge a defensive decision boundary against OoD samples via fine-tuning. Unlike classification models, object detectors are inherently trained to differentiate foreground objects from background objects, often using the objectness score. As OoD filtering is only triggered when the object detector acknowledges the existence of an object, reducing the false detection can be done as a joint approach: In addition to equipping a filter similar to that of VOS or BAM, additionally adjust the decision boundary of the object detector to reduce opportunities to trigger false detections. Specifically, we introduce a fine-tuning strategy that uses a set of OoD samples proximal to ID data - termed “proximal OoD” - which can be inferred from the ID data distribution without relying on specific test-time OoD datasets. From the perspective of learned “concepts”, these samples share high-level semantic features with ID data while differing in low-level attributes, effectively bridging the gap between ID and truly novel OoD instances. During fine-tuning of the models on these proximal OoD samples, the model is trained to treat proximal OoD samples as background, thereby shaping an implicit defensive decision boundary against general OoD instances. Our experimental results show that this method effectively mitigates hallucinations when encountering both near-OoD and far-OoD inputs, demonstrating its efficacy.
@@
  • We identify key dataset quality issues in existing OoD benchmarks that negatively impact the evaluation of OoD detection methods in YOLO models.
@@
  • We conduct extensive experiments that demonstrate the effectiveness of our approach, achieving substantial improvements in both near-OoD and far-OoD detection.
@@ II RELATED WORK
Overconfidence or hallucination is a symptom that can arise in both ID and OoD scenarios, where our work focuses on the OoD case. OoD detection has gained significant attention in computer vision. Within image classification tasks, techniques leveraging confidence scores [7, 8, 9], feature-space distances [10, 11, 12, 13] or generative modeling [14, 15] can detect OoD anomalies. We refer readers to a recent survey paper [16] for a detailed summary of techniques. Due to the maturity of OoD detection research in classification, the community has developed a standard benchmark OpenOoD [17], enabling researchers to compare against different techniques. For OoD detection in object detection, the problem is less explored despite its importance, and as of the time of writing (March 2025), OpenOoD has not yet supported object detection. Methods operated on Faster-RCNN, including VOS [2], SAFE [3] and BAM [4] enable reducing hallucination in OoD samples via refining feature space representations and uncertainty estimation. The recent study [6] translated multiple OoD detection methods within classification to the YOLO-family object detection settings, including maximum logits (MLS) [18], maximum softmax probability (MSP) [19], energy-Based methods (EBO) [8], Mahalanobis distance (MDS) [10], k-nearest neighbors (KNN) [12] and the recent SCALE method [9]. While the performance of reducing hallucination against OoD samples has been considerably superior to those applied on Faster-RCNN, it is still far from satisfactory. Our result in highlighting problems within the existing object detection OoD benchmark dataset implies that the performance of prior results should be better than the reported numbers. It also stresses the importance of having a rigorous benchmark to ensure an unbiased performance evaluation. Furthermore, our proposed fine-tuning based on introducing proximal OoD examples in the detector is experimentally proven to be highly effective. It also reflects the design principle that reducing hallucination when encountering novel or abnormal objects should be a joint effort (similar to adversarial training [20] for adversarial robustness) between the object detector and the subsequent filter that performs OoD detection.
@@ III-A2 Type 1 data error: ID Objects in OoD Test Sets
We can now describe the Type 1 data error and its associated consequences. A Type 1 error occurs when an image in the OoD-only test set contains at least one ID object (i.e., belonging to $\mathcal{O}_{\text{in}}$). An example of this issue is shown in Fig. 1(a). Precisely, an OoD-only test set $\mathcal{D}_{\text{test}}^{\text{OoD}}$ manifests Type 1 data errors iff the following holds:
@@ III-A3 How ID-only data set is used for deciding FPR95 decision boundary for the OoD filter
provided $\;\;\;\;\;\text{provided that}\;\;I\in\mathcal{D}_{\text{cali}}^{\text{ID}}$
@@ III-A4 Type 2 data issue: Unlabelled OoD Objects in ID-only data sets
Unfortunately, the above experiment design is based on an assumption where there exist no OoD objects that an object detector $f$ can be wrongly triggered. Precisely, an ID-only dataset $\mathcal{D}_{\text{cali}}^{\text{ID}}$ manifests Type 2 data issues iff the following holds, where $\hat{y}_{i}$ is the associated ground truth:
@@ III-C ID Outlier-induced Ambiguity
III-D New Benchmarks
As discussed in previous sections, existing OoD benchmark datasets are prone to dual mislabeling errors, limiting their reliability for OoD evaluation. We introduce two new OoD benchmark datasets, Near-OoD and Far-OoD  [^2] https://gricad-gitlab.univ-grenoble-alpes.fr/dnn-safety/m-hood, designed based on image semantics and empirical difficulty to address this issue. Near-OoD includes objects semantically similar to training categories, often presenting empirical challenges for the model due to their resemblance to ID classes. At the same time, Far-OoD consists of more semantically distant objects, including background images without recognizable objects. Each dataset consists of approximately 1,000 images sourced from the OpenImagesV7 dataset, covering over 500 diverse object detection categories. To ensure the integrity of these benchmarks, we verify that none of the images contain objects from the training categories. This is initially done by performing inference with a pre-trained, high-accuracy YOLO model on OpenImagesV7. We then analyze the model’s predictions and filter out images containing indisputable objects from training categories. Finally, a manual inspection is conducted for further validation.
@@ IV FINE-TUNING FOR MITIGATING HALLUCINATION
This section introduces an alternative strategy to mitigate YOLO’s hallucination (overly confident predictions) on OoD samples without explicitly performing OoD detection. Our approach leverages proximal OoD samples, which are semantically similar to the ID categories but do not belong to them. This fine-tuning method, similar to adversarial training, adjusts the model’s decision boundaries to make it more conservative when facing OoD samples. This reduces the likelihood of hallucinated predictions, whereas adversarial training focuses on enhancing the model’s robustness against adversarial examples. Rather than training a model from scratch, we adopt a fine-tuning strategy that is computationally efficient and effective in reshaping the decision boundary. By guiding the model to differentiate between proximal OoD data and ID samples, we hypothesize that it can better generalize and distinguish broader OoD instances. We first describe the construction of the proximal OoD dataset and then introduce our fine-tuning methodology.
@@ IV-A Proximal OoD Dataset Construction
To construct a high-quality proximal OoD dataset, we leverage Caltech 256 [25], a diverse object-centric dataset, as a source of candidate samples. Given the large number of categories, manual selection is impractical. Instead, we employ Deepseek-R1, a reasoning-based large language model, to generate candidate categories that align with our selection principles, followed by manual verification to ensure their validity. Fig. 5 presents the prompt template used to generate candidate proximal OoD categories. Once the proximal OoD categories are identified, we collect approximately 2,000 images and ensure they do not contain any ID objects, as they will serve as background or negative samples during fine-tuning. We systematically analyze the dataset using a YOLOv8 model pre-trained on OpenImagesV7 dataset to enforce this constraint. Specifically, we run inference on the collected images and use the FiftyOne library to filter out images containing ID objects based on the inference results. This automated refinement minimizes ID contamination while preserving the diversity of OoD samples. Finally, the curated dataset is converted into the YOLO format, ensuring seamless integration into the fine-tuning pipeline. This approach streamlines dataset construction and enhances the reliability of proximal OoD selection, enabling more effective mitigation of overconfident predictions.
@@ V-A Experimental Setup
We evaluate our method on two well-trained YOLOv10 models from [26], trained on PASCAL-VOC and BDD-100K, respectively. We use the calibrated near-OoD and far-OoD datasets introduced in the previous subsection to assess hallucinations. The effectiveness of fine-tuning is measured by its ability to reduce hallucinations, which we quantify using OoD counts—the number of predictions on OoD data. Ideally, this value should be zero. However, ensuring that the model’s primary detection performance is not compromised beyond mitigating hallucinations is crucial. To this end, we track mean average precision (mAP) to evaluate the model’s accuracy on ID data.

### v3

# Revisiting Out-of-Distribution Detection in Real-time Object Detection: From Benchmark Pitfalls to a New Mitigation Paradigm

@@ Abstract
Out-of-distribution (OoD) inputs pose a persistent challenge to deep learning models, often triggering overconfident predictions on non-target objects. While prior work has primarily focused on refining scoring functions and adjusting test-time thresholds, such algorithmic improvements offer only incremental gains. We argue that a rethinking of the entire development lifecycle is needed to mitigate these risks effectively. This work addresses two overlooked dimensions of OoD detection in object detection. First, we reveal fundamental flaws in widely used evaluation benchmarks: contrary to their design intent, up to 13% of objects in the OoD test sets actually belong to in-distribution classes, and vice versa. These quality issues severely distort the reported performance of existing methods and contribute to their high false positive rates. Second, we introduce a novel training-time mitigation paradigm that operates independently of external OoD detectors. Instead of relying solely on post-hoc scoring, we fine-tune the detector using a carefully synthesized OoD dataset that semantically resembles in-distribution objects. This process shapes a defensive decision boundary by suppressing objectness on OoD objects, leading to a 91% reduction in hallucination error of a YOLO model on BDD-100K. Our methodology generalizes across detection paradigms such as YOLO, Faster R-CNN, and RT-DETR, and supports few-shot adaptation. Together, these contributions offer a principled and effective way to reduce OoD-induced hallucination in object detectors. Code and data are available at: https://gricad-gitlab.univ-grenoble-alpes.fr/dnn-safety/m-hood.
Index Terms:
Object detection, Out-of-distribution detection, Benchmark calibration, New mitigation strategy, Ojectness-guided fine-tuning, YOLOs, Faster-RCNN, RT-DETR.
@@ I INTRODUCTION
Deep learning models, such as image classifiers and object detectors, have achieved remarkable success in accomplishing in-distribution (ID) tasks by learning meaningful patterns. However, these models tend to overgeneralize when exposed to out-of-distribution (OoD) inputs, i.e., samples that fall outside the training distribution and do not correspond to any ID category. Predictions on OoD inputs often result in confident yet nonsensical or irrelevant outputs. We refer to this failure mode as hallucination in computer vision. The risks associated with such OoD-induced hallucinations are twofold. They may arise unintentionally, as in type errors or semantic bugs in code, where the syntax is correct but the meaning is wrong. Alternatively, they can be exploited deliberately, allowing adversaries to induce high-confidence false detections without requiring perturbations or access to training data.
@@
Figure 2: Reduction of hallucination errors on OoD datasets. Our approach significantly reduces hallucinated detections on OoD inputs across two ID datasets (BDD100K and PASCAL-VOC) and multiple detection architectures (YOLO, Faster R-CNN, RT-DETR). The method achieves substantial reductions in spurious detections—up to 91%—and generalizes well across different detector families, including one-stage, two-stage, and Transformer-based models.
To mitigate the risks posed by such illegal inputs, OoD detection has emerged as a promising solution, attracting extensive research efforts focused on designing effective test-time filtering algorithms. For instance, in image classification, these methods typically assign low confidence scores to OoD samples and redirect their predictions to a dedicated “OoD” class. Inspired by this paradigm, recent techniques such as VOS [1], SAFE [2], and BAM [3] are implemented on top of Faster-RCNN [4] as an extra component to subsequently filter the detected and classified objects by setting these objects to be of class “OoD”.
While prior work has largely focused on refining OoD scoring functions and adjusting test-time thresholds, we argue that such algorithmic efforts for OoD detection, though valuable, offer only incremental progress in the context of object detection. Unlike classification tasks, which produce a single prediction per image, object detection involves structured outputs with multiple objects and spatial reasoning, making the detection and mitigation of OoD instances considerably more complex. A recent study [5] illustrates this challenge by adapting leading classification-based OoD techniques to YOLO detectors, only to find a substantial drop in performance compared to their application on Faster R-CNN [3]. These findings suggest that post-hoc algorithms alone are insufficient. We contend that optimizing OoD detection performance cannot rely on a single algorithmic solution. Instead, it calls for a re-examination of the entire development lifecycle of current mitigation paradigms in object detection. In this work, we address two underexplored aspects: (1) fundamental implementation flaws in existing OoD detection benchmarks, and (2) the lack of training-time mitigation strategies that explicitly leverage the structural properties of object detectors, independent of external OoD filters.
First, we identify dataset quality issues and benchmark design in evaluating OoD used in VOS [1] and all other techniques, contributing to the high false positive rates observed in previous studies. Specifically, we uncover two critical problems and study their impact both in theory and concrete empirical evaluations:
@@
Second, to reduce hallucinations (overconfident predictions) caused by OoD samples, we propose a new mitigation paradigm that forges a defensive decision boundary against OoD inputs through fine-tuning. Unlike classification models, object detectors are inherently trained to differentiate foreground objects from background objects, often using the objectness score. As OoD filtering is only triggered when the object detector acknowledges the existence of an object, reducing the false detection can be done as a joint approach: In addition to equipping a filter similar to that of VOS or BAM, additionally adjust the decision boundary of the object detector to reduce opportunities to trigger false detections. Specifically, we introduce a fine-tuning strategy that makes use of a set of OoD samples proximal to ID data - termed “proximal OoD” - which can be inferred from the ID data distribution without relying on specific test-time OoD datasets. From the perspective of learned “concepts”, these samples share high-level semantic features with ID data while differing in low-level attributes, effectively bridging the gap between ID and truly novel OoD instances. During fine-tuning of the models on these proximal OoD samples, the model is trained to treat proximal OoD samples as background, thereby shaping an implicit defensive decision boundary against general OoD instances. Our experimental results, as shown in Fig. 2, demonstrate that the proposed method effectively mitigates hallucinations across different experimental settings when encountering both near-OoD and far-OoD inputs, demonstrating its efficacy.
@@
  • We identify critical dataset quality issues in existing OoD benchmarks and develop an automated tool for constructing customized evaluation datasets tailored to an object detector’s training categories.
@@
  • We show through large-scale experiments that our approach generalizes effectively across diverse detection paradigms, including single-stage CNN-based models (e.g., YOLO [6]), two-stage models (e.g., Faster R-CNN [4]), transformer-based detectors (e.g., RT-DETR [7]), and few-shot fine-tuning settings.
  • We provide additional insights into the effectiveness of our method through explainability and confidence shift analysis.
The remainder of this paper is organized as follows: Section II reviews related work in OoD detection for image classification and object detection. Section III details our analysis of dataset quality issues. Section IV presents our proposed fine-tuning approach. Section V reports the experimental results, and Section VI analyzes the effectiveness of our approach. Section VII discusses the open challenges in mitigating OoD inputs for object detection. Section VIII concludes the paper with discussions on future directions.
@@ II RELATED WORK
Figure 3: Positioning of our training-time OoD safety method (green) within the broader defense landscape. Adversarial defenses (red) improve robustness against malicious perturbations through adversarial training and inference-time defenses. In parallel, our method (green) proactively shapes decision boundaries to handle risky OoD inputs. This previously overlooked training-time safety mechanism complements post hoc filtering and aligns with a broader AI safety paradigm: combining robustness and reliability across the training–inference pipeline.
@@
Overconfidence or hallucination is a symptom that can arise in both ID and OoD scenarios, where our work focuses on the OoD case. OoD detection has gained significant attention in computer vision. Within image classification tasks, techniques leveraging confidence scores [8, 9, 10], feature-space distances [11, 12, 13, 14] or generative modeling [15, 16] can detect OoD anomalies. We refer readers to a recent survey paper [17] for a detailed summary of techniques. Due to the maturity of OoD detection research in classification, the community has developed a standard benchmark OpenOoD [18], enabling researchers to compare against different techniques. For OoD detection in object detection, the problem is less explored despite its importance, and as of the time of writing (March 2025), OpenOoD has not yet supported object detection. Methods operated on Faster-RCNN, including VOS [1], SAFE [2] and BAM [3] enable reducing hallucination in OoD samples via refining feature space representations and uncertainty estimation. The recent study [5] translated multiple OoD detection methods within classification to the YOLO-family object detection settings, including maximum logits (MLS) [19], maximum softmax probability (MSP) [20], energy-based methods (EBO) [9], Mahalanobis distance (MDS) [11], k-nearest neighbors (KNN) [13] and the recent SCALE method [10]. While the performance of reducing hallucination against OoD samples has been considerably superior to those applied on Faster-RCNN, it is still far from satisfactory. Our result in highlighting problems within the existing object detection OoD benchmark dataset implies that the performance of prior results should be better than the reported numbers. It also stresses the importance of having a rigorous benchmark to ensure an unbiased performance evaluation.
Furthermore, our fine-tuning strategy, based on injecting proximal OoD examples during training, marks a shift in methodology within the realm of OoD detection for object detectors. Prior efforts have largely focused on designing inference-time filters to reject OoD objects. In contrast, our approach leverages the unique structure of object detectors—particularly the objectness score—to proactively shape safer decision boundaries during training. This constitutes a training-time safety mechanism that complements traditional post hoc filtering. Such a perspective aligns with a broader trend in machine learning safety: just as adversarial robustness [21] research has evolved to include both adversarial training and inference-time defenses, we argue that OoD safety should be treated as a joint endeavor across the training–inference pipeline. As illustrated in Fig. 3, adversarial training improves resilience against malicious, fine-grained perturbations, while our method improves reliability under semantically incompatible inputs. Together, these complementary defenses help characterize a more holistic safety framework for open-world object detection.
This work builds upon our earlier conference version [22], which focused on YOLO-based object detectors. We substantially extend that study by (i) developing an automated OoD benchmarking tool, (ii) generalizing to multiple detection paradigms including Faster R-CNN and RT-DETR, (iii) exploring few-shot fine-tuning strategies, and (iv) providing in-depth analyses of our method’s effectiveness.
@@ III-A2 Type 1 data error: ID Objects in OoD Test Sets
We can now describe the Type 1 data error and its associated consequences. A Type 1 error occurs when an image in the OoD-only test set contains at least one ID object (i.e., belonging to $\mathcal{O}_{\text{in}}$). An example of this issue is shown in Fig. 1(a). Precisely, an OoD-only test set $\mathcal{D}_{\text{test}}^{\text{OoD}}$ manifests Type 1 data errors if and only if the following holds:
@@ III-A3 How ID-only data set is used for deciding FPR95 decision boundary for the OoD filter
$\;\;\;\;\;\text{provided that}\;\;I\in\mathcal{D}_{\text{cali}}^{\text{ID}}$
@@ III-A4 Type 2 data issue: Unlabelled OoD Objects in ID-only data sets
Unfortunately, the above experiment design is based on an assumption where there exist no OoD objects that an object detector $f$ can be wrongly triggered. Precisely, an ID-only dataset $\mathcal{D}_{\text{cali}}^{\text{ID}}$ manifests Type 2 data issues if and only if the following holds, where $\hat{y}_{i}$ is the associated ground truth:
@@ III-C ID Outlier-induced Ambiguity
III-D Automated Data Curation Pipeline
Both our benchmarking and fine-tuning strategies depend on high-quality, task-specific image data. However, manually collecting such data is labor-intensive—particularly when adapting to diverse sources or meeting the specific demands of evaluation and fine-tuning. To address this challenge, we introduce a universal automated data curation pipeline that streamlines the construction of customized datasets across different experimental stages. This pipeline integrates the FiftyOne library [27] with the YOLOE  [28] open-vocabulary object detection model to systematically gather and curate relevant images while ensuring the exclusion of ID content.
As illustrated in Fig. 7, the pipeline workflow begins with users defining task-specific categories tailored to their requirements, whether for fine-tuning, testing, or few-shot learning scenarios. Leveraging the FiftyOne Dataset Zoo API, it then retrieves candidate images for these categories from selected data sources, with configurable per-category image counts to accommodate various experimental needs. The automated filtering process employs the pre-trained YOLOE model, configured with prompts corresponding to the predefined categories. To maintain dataset purity, the pipeline retains only images that are free from predicted or annotated ID objects. For additional quality assurance, an optional human-in-the-loop verification step is available through FiftyOne’s interactive visualization and annotation tools. The final curated dataset is exported in standard format, ensuring seamless integration with downstream training and evaluation pipelines. This automated and modular approach streamlines the construction of high-quality object detection datasets, maintaining consistency and reproducibility regardless of the underlying data source or specific task requirements.
Figure 7: Overview of our universal pipeline for automated data curation.
III-E New Benchmarks
To address the limitations identified in existing OoD benchmarks (see previous sections), we leverage our automated data curation pipeline to construct two new evaluation datasets: Near-OoD and Far-OoD  [^2] https://huggingface.co/datasets/HugoHE/m-hood-dataset, specifically designed to enable fine-grained analysis of hallucination suppression. Near-OoD comprises images containing objects that are visually and semantically similar to the training categories, presenting a particularly challenging scenario for object detectors. In contrast, Far-OoD contains images with objects that are distinctly different from the training set, as well as backgrounds without recognizable objects. Each benchmark contains 1,000 images, carefully sampled from over 500 diverse categories in OpenImages V7. These new benchmark datasets allow for a more reliable and challenging evaluation for OoD detection for object detection models.
@@ IV FINE-TUNING FOR MITIGATING HALLUCINATION
This section introduces an alternative strategy to mitigate YOLO’s hallucination (overly confident predictions) on OoD samples without explicitly performing OoD detection. Our approach leverages proximal OoD samples, which are semantically similar to the ID categories but do not belong to them. This fine-tuning method, illustrated in Fig. 8, similar to adversarial training, adjusts the model’s decision boundaries to make it more conservative when facing OoD samples. This reduces the likelihood of hallucinated predictions, whereas adversarial training focuses on enhancing the model’s robustness against adversarial examples. Rather than training a model from scratch, we adopt a fine-tuning strategy that is computationally efficient and effective in reshaping the decision boundary. By guiding the model to differentiate between proximal OoD data and ID samples, we hypothesize that it can better generalize and distinguish broader OoD instances. We first describe the construction of the proximal OoD dataset and then introduce our fine-tuning methodology.
Figure 8: Conceptual two-stage decision-making logic in object detection. Regardless of architectural design (e.g., YOLO, Faster R-CNN, RT-DETR), modern detectors follow a common logical process: first, candidate regions are scored by objectness to filter out background or ambiguous inputs; second, high-scoring regions are classified and localized. Our method fine-tunes the objectness scoring mechanism using proximal OoD data, encouraging low objectness scores for OoD inputs and enabling early suppression prior to class prediction.
@@ IV-A Proximal OoD Dataset Construction
After selecting suitable proxy categories, we use our automated data curation pipeline (see Section III-D) to extract approximately 2,000 images from the Caltech 256 dataset as proximal OoD samples for fine-tuning. Caltech 256 provides a broad and diverse set of object-centric images, making it particularly suitable for capturing the subtle semantic proximity to in-distribution classes. Importantly, sourcing data from Caltech 256 ensures no overlap or leakage with the evaluation benchmarks, thereby preserving the integrity of our experiments.
@@ V-A Experimental Setup
We evaluate our method on two well-trained YOLOv10 models from [6], trained on PASCAL-VOC and BDD-100K, respectively. We use the calibrated Near-OoD and Far-OoD datasets introduced in the previous subsection to assess hallucinations. We use a confidence score threshold of 0.25 for all models, following the default setting in the YOLO series models [29]. The effectiveness of fine-tuning is measured by its ability to reduce hallucinations, which we quantify using OoD counts—the number of predictions on OoD data. Ideally, this value should be zero. However, ensuring that the model’s primary detection performance is not compromised beyond mitigating hallucinations is crucial. To this end, we track mean average precision (mAP) to evaluate the model’s accuracy on ID data.